The profit made by a company when 60 units of its product is sold is R 1 600.00. When 150 units of its products are sold, the profit increases to R 5 200.00. Assuming that the profit function is linear and of the form
1
Expert's answer
2019-04-03T10:07:21-0400

We have two points with coordinates: A(60, 1600) and B(150, 5200). Since the function is linear, we use the formula "\frac {x - x_1} {x_2-x_1}=\frac {y-y_1} {y_2-y_1}" (

"x_1=60"

"x_2=150"

"y_1=1600"

"y_2=5200"


"\frac {x-60} {90} = \frac {y-1600} {3600}"

"(x-60)*3600=(y-1600)*90"

"90*y=3600*x-72000"

"y=40*x-800"